Caspian Sea Drinks is considering the purchase of a new water filtration system produced by Rube Goldberg Machines. This new equipment, the RGM-7000, will allow Caspian Sea Drinks to expand production. It will cost $12.00 million fully installed and will be fully depreciated over a 15 year life, then removed for no cost. The RGM-7000 will result in additional revenues of $2.53 million per year and increased operating costs of $621,174.00 per year. Caspian Sea Drinks' marginal tax rate is 22.00%. The internal rate of return for the RGM-7000 is _____.
Answer #1

A B с D 1 2 Particulars Inflow(Years 1-15) 3 Revenue 25,30,000.00 4 Expense 6,21,174.00 5 EBITDA 19,08,826.00 =Revenue-Expens

A B 1 Inflow(Years 1-15) 2530000 621174 =B3-B4 =Revenue-Expense =12000000/15 =Cost of asset/number of years =B5-B7 =B9*0.22 =